A 9.00 kg mass is moving to the right with a velocity of 14.0 m/s. A 12.0 kg mass is moving to the left with a velocity of 5.00 m/s. Assuming that these two balls have a head on collision and stick together, what will be the final velocity of the combination?


1
Expert's answer
2021-04-12T07:00:42-0400
"MV-mv=(m+M)u\\(9)(14)-(5)(12)=(9+12)u\\u=3.14\frac{m}{s}"